OpenStudy (anonymous):

Which point lies on a circle that is centered at A(3, 3) and passes through B(6, 5)?

OpenStudy (solomonzelman):

you know the center. so you know the \(\Large\color{blue}{ \underline{(x-h)^2+(y-k)^2}=r }\) To find the radius, find the distance between the 2 points.

OpenStudy (anonymous):

what do i plug in for h and k?

OpenStudy (anonymous):

(h,k) is the centre

OpenStudy (solomonzelman):

(h,k) is the center of the circle. And you are given that the center is (3,3)

OpenStudy (anonymous):

ohh okay so i just plug it in then its like 13?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ (x-\underline{h})^2+(y-\underline{k})^2=r }\) \(\Large\color{blue}{ (x-\underline{3})^2+(y-\underline{3})^2=r }\)

OpenStudy (solomonzelman):

Now, find the distance. The distance between (x1,y1) and (x2,y2) is given by the formula, \(\LARGE\color{black}{ distance=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2} }\)

OpenStudy (anonymous):

3.6?

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ distance=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2} }\) \(\LARGE\color{black}{ distance=\sqrt{(3-5)^2+(3-6)^2} }\) \(\LARGE\color{black}{ distance=\sqrt{(-2)^2+(-3)^2} }\) \(\LARGE\color{black}{ distance=\sqrt{4+9} }\) \(\LARGE\color{black}{ distance=\sqrt{11} }\) \(\LARGE\color{black}{ distance≈3.3166 }\)

OpenStudy (solomonzelman):

about 3.32

OpenStudy (anonymous):

4+9 isnt eleven though

OpenStudy (solomonzelman):

And the distance between the 2 point is/was the radius. So how would you write the equation of the circle ?